In risk management, insurance is an example of:

Select one:

a. Risk avoidance

b. Risk deflection

c. Risk acceptance

d. Contingency planning

There is a definite link between the level of risk an enterprise encounters and project time frame.

Select one:

a. true

b. false

If review of a project's status indicates that EV = $400, AC = $400, and PV = $500, the project is:

Select one:

a. On budget, behind schedule

b. On budget, ahead of schedule

c. Over budget, behind schedule

d. Over budget, ahead of schedule

The bottom-most level of the WBS -- the level at which project budget and schedule data are captured -- is called:

Select one:

a. Work package

b. Code of accounts

c. Budget baseline

d. Change control level
Answer #1


In risk management, insurance is an example of

b.Risk deflection

Insurance is a way of transferring the risk i.e risk deflection

There is a definite link between the level of risk an enterprise encounters and project time frame.

Correct answer is True

Longer the project duration, higher chance of risk affecting project

If review of a project's status indicates that EV = $400, AC = $400, and PV = $500, the project is:

SPI = EV/PV = 400/500 = 0.8

SPI <1 suggests that project is behind schedule

CPI = EV/AC = 400/400 = 1

CPI = 1 suggests that project is on budget

Correct answer is a. On budget, behind schedule


The bottom-most level of the WBS -- the level at which project budget and schedule data are captured -- is called:

correct answer is a. Work package